OpenStudy (anonymous):

Hey guys i need your help without graphing, determine whether the function represents exponential growth or decay. y=3(1.04)^x y=2(0.95)^x y=2(5/3)^x thanks

jimthompson5910 (jim_thompson5910):

In general, exponential functions can be written in the form \[\Large y = a*(b)^x\] the term 'a' is the starting amount (when x = 0) The b tells you if you have exponential growth or decay If b < 1, then you have exponential decay If b > 1, then you have exponential growth

OpenStudy (anonymous):

so the equation y=3(1.04)^x would be decay?

jimthompson5910 (jim_thompson5910):

what are a and b in the equation y=3(1.04)^x

OpenStudy (anonymous):

a=3 and b=1.04

jimthompson5910 (jim_thompson5910):

now look at the rules I posted for exponential decay and exponential growth

OpenStudy (anonymous):

oh that's right thank you so much

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

just to make sure here what i got y=3(1.04)^x would be decay y=2(0.95)^x= would be growth y=2(5/3)x= would be decay if it's wrong can you corrected

jimthompson5910 (jim_thompson5910):

you have it all backwards

jimthompson5910 (jim_thompson5910):

if b = 1.04, then b > 1, so we have growth (not decay)

OpenStudy (anonymous):

ops thank you very much